OpenStudy (anonymous):

How do you take the derivative of x raised to a complex number? For example, let z = a + b*i and let y = x^z. How do you get dy/dx? I assume that you'd bring down the (a + b*i) to multiply the x, but what is the proper way to decrement the exponential term?

OpenStudy (anonymous):

Take ln!

OpenStudy (anonymous):

So long as b is zero, you multiply the exponent by a and subtract one from the exponent\[ax ^{a-1}\]I don't know about other cases of b

OpenStudy (anonymous):

Ln y = ( a + bi) Ln x -> y'/y = ( a + bi) / x => y' = ( a + bi) y/x

OpenStudy (anonymous):

Thanks, how did you get the 2nd step?

OpenStudy (anonymous):

Thanks @Chlorophyll !

OpenStudy (anonymous):

The way you write it, the power rule is true for all numbers. So, when solving your equation @schmidtsl just use the power rule I wrote above.

OpenStudy (anonymous):

Yes, I see that Chlorophyll's result leads to exactly that (which is what you'd expect). Thanks.

OpenStudy (anonymous):

Oh, now I see what Chlorophyll meant. In the 2nd step, you take the derivative with respect to x on both sides. d(ln(y))/dx = (1/y) * dy/dx = y'/y and d(ln(x))/dx = 1/x. Thank you!!

OpenStudy (anonymous):

@L.T. @schmidtsl Sorry, I was away from my PC!